OpenStudy (anonymous):

When 2 moles of KF are placed in one liter of water, what happens to the boiling point? a. it decrease by 3.72 degree Celsius b. it decreases by 7.44 degrees Celsius c. it increases by 2.08 degrees Celsius d. it increases by 1.04 degrees Celsius The answer is C but I don't know how to actually solve it... Please help.

OpenStudy (aaronq):

use the formula: \(\Delta T=i*m*K_{\theta}\) where, \(\theta=b\) when finding BP, and \(\theta=f\) when finding freezing point (By that i mean use the appropriate constant.) More on this: http://openstudy.com/study#/updates/521ecb3ae4b0750826e0c362

OpenStudy (anonymous):

@aarong I am very lost on this subject....

OpenStudy (eric_d):

@abb0t

OpenStudy (eric_d):

I think.. v got to find 'i' first frm KF

OpenStudy (anonymous):

Ok what formula is this and what's I supposed to be?

OpenStudy (anonymous):

@matt101 please can you help?

OpenStudy (eric_d):

KF

OpenStudy (eric_d):